# MoRAG - Multi-Fusion Retrieval Augmented Generation for Human Motion

### v1

@@ Abstract
We introduce MoRAG, a novel multi-part fusion based retrieval-augmented generation strategy for text-based human motion generation. The method enhances motion diffusion models by leveraging additional knowledge obtained through an improved motion retrieval process. By effectively prompting large language models (LLMs), we address spelling errors and rephrasing issues in motion retrieval. Our approach utilizes a multi-part retrieval strategy to improve the generalizability of motion retrieval across the language space. We create diverse samples through the spatial composition of the retrieved motions. Furthermore, by utilizing low-level, part-specific motion information, we can construct motion samples for unseen text descriptions. Our experiments demonstrate that our framework can serve as a plug-and-play module, improving the performance of motion diffusion models. Code, pretrained models and sample videos will be made available at: https://motion-rag.github.io/
@@ 1 Introduction
Table 1: Comparision of text-to-motion retrieval approaches
@@ 3.2 Generation of part-specific descriptions
Specifically prompting for the position provides the global orientation of the body parts which results in better retrieval.
@@ 3.3 Multi-part motion retrieval
As shown in Fig. 4, MoRAG uses $3$ independently trained TMR[23] models, $MoRAG=\{MoRAG_{torso},MoRAG_{hands},MoRAG_{legs}\}$ corresponding to the respective body part. We do not train separate left and right body parts models to avoid asynchronous movements in the composed motion. For a part $p\in\{torso,hands,legs\}$, the retrieval model $MoRAG_{p}=\{T^{Enc}_{p},M^{Enc}_{p},M^{Dec}_{p},\mathcal{D}^{Enc}_{p}\}$ consists of a text encoder, motion encoder, motion decoder, and motion database respectively. The model architecture for the encoder and decoder are based on TEMOS[22].
@@
To enhance the effectiveness of motion retrieval, we train all three motion encoders ($M^{Enc}_{p}$) using full-body motion sequences instead of just the respective body parts. This approach is based on the observation that LLM-generated part-specific descriptions contain information about the queried body part about other body parts. Utilizing full-body motion sequences allows us to leverage intra-joint information, resulting in more coherent and semantically accurate motion retrieval.
@@
3.4 Spatial motion composition
The retrieved motion sequences {$R^{1}_{p},\dots,R^{i}_{p},\dots,R^{k}_{p}$} are composed such that the $i_{th}$ sequence corresponding to each part $p$ is used to construct the $i_{th}$ full-body motion sequence $C^{i}$. This results in $k$ full-body motion sequences {$C^{1},\dots,C^{i},\dots,C^{k}$}, which are used as additional guidance for the motion diffusion model. We followed a rank-by-rank combination approach to generate these top-$k$ sequences. However, alternative combination methods could be employed to create a significantly larger number of sequences. Our composition approach is similar to SINC[4] but we do not require the use of an LLM for mapping joints from the retrieved sequences to the composed sequence.
$C^{i}_{f}[j_{p}]=R^{i}_{p,f}[j_{p}]$ (9)
$p\in\{torso,hands,legs\},f\in[1,f_{min}]$
@@ 4 Experiments & Results
First, we describe the dataset (Sec. 4.1) and implementation details (Sec.4.2) used in our experiments. Following that, we analyze the effectiveness of the MoRAG framework, comparing it to previous research works by providing an analysis of both retrieval and generation results. (Sec.4.3).
4.1 Dataset
4.2 Implementation Details
We use OpenAI’s GPT-3.5-turbo-instruct as the large language model (LLM) due to its efficiency in understanding and executing specific instructions and its ability to provide direct answers to questions. The proposed prompting strategy consumes a maximum of 256 tokens together for prompt and generation. We use the completions API endpoint with the default parameters to generate the desired part-specific descriptions.
@@ 5 Conclusion
Future work could extend our approach to other architecture-based generative models. Incorporating additional body part information, such as fingers, head, and lip movements from respective part-specific databases, would enhance the realism of generated samples and better handle more complex language descriptions. Furthermore, our approach can be applied to create new data samples, useful for both training and guiding motion generation models, thereby expanding its potential to handle unusual inputs.

### v2

@@ Abstract
We introduce MoRAG, a novel multi-part fusion based retrieval-augmented generation strategy for text-based human motion generation. The method enhances motion diffusion models by leveraging additional knowledge obtained through an improved motion retrieval process. By effectively prompting large language models (LLMs), we address spelling errors and rephrasing issues in motion retrieval. Our approach utilizes a multi-part retrieval strategy to improve the generalizability of motion retrieval across the language space. We create diverse samples through the spatial composition of the retrieved motions. Furthermore, by utilizing low-level, part-specific motion information, we can construct motion samples for unseen text descriptions. Our experiments demonstrate that our framework can serve as a plug-and-play module, improving the performance of motion diffusion models. Code, pre-trained models, and sample videos are available at motion-rag.github.io.
@@ 1 Introduction
Table 1: Comparision of text-to-motion retrieval approaches - Text Robustness (ability to handle diverse language inputs), Generalizability (adaptation to similar yet altered data), Diversity (capacity to produce varied outputs), and Zero-Shot Setting (performance on previously unseen data).
@@ 3.2 Generation of part-specific descriptions
Specifically prompting for the position provides the global orientation of the body parts which results in better retrieval. Further details along with rendered video showcasing significance of position information can be found in project page.
@@ 3.3 Multi-part motion retrieval
As shown in Fig. 4, MoRAG uses $3$ independently trained TMR[23] models, $MoRAG=\{MoRAG_{torso},MoRAG_{hands},MoRAG_{legs}\}$ corresponding to the respective body part. We do not train separate left and right body parts models to avoid asynchronous movements in the composed motion. Rendered video showcasing the shortcomings of seperate left and right body parts retrieval can be found in project page. For a part $p\in\{torso,hands,legs\}$, the retrieval model $MoRAG_{p}=\{T^{Enc}_{p},M^{Enc}_{p},M^{Dec}_{p},\mathcal{D}^{Enc}_{p}\}$ consists of a text encoder, motion encoder, motion decoder, and motion database respectively. The model architecture for the encoder and decoder are based on TEMOS[22].
@@
To enhance the effectiveness of motion retrieval, we train all three motion encoders ($M^{Enc}_{p}$) using full-body motion sequences instead of just the respective body parts. This approach is based on the observation that LLM-generated part-specific descriptions contain information about the queried body part about other body parts. Utilizing full-body motion sequences allows us to leverage intra-joint information, resulting in more coherent and semantically accurate motion retrieval. LLM-generated part-specific descriptions for sample text inputs alongside their corresponding retrieval results can be found in project page.
@@
3.4 Spatial motion composition
The retrieved motion sequences {$R^{1}_{p},\dots,R^{i}_{p},\dots,R^{k}_{p}$} are composed such that the $i_{th}$ sequence corresponding to each part $p$ is used to construct the $i_{th}$ full-body motion sequence $C^{i}$. This results in $k$ full-body motion sequences {$C^{1},\dots,C^{i},\dots,C^{k}$}, which are used as additional guidance for the motion diffusion model. To generate these top-$k$ sequences, we employed a rank-by-rank combination approach, where the retrieved part-specific sequences are selected and combined in order of their rank across different models. However, alternative combination methods could be employed to create a significantly larger number of sequences. Our composition approach is similar to SINC[4] but we do not require the use of an LLM for mapping joints from the retrieved sequences to the composed sequence.
$C^{i}_{f}[j_{p}]=R^{i}_{p,f}[j_{p}]$ (9)
$p\in\{torso,hands,legs\},f\in[1,f_{min}]$
@@ 4 Experiments & Results
4.1 Dataset and Implementation Details
We chose HumanML3D[12] to evaluate our framework due to its extensive and diverse collection of text-annotated motions. It provides annotations for the motions in the AMASS[21] and HumanAct12[13] datasets. We follow the same splits as provided in TMR[23] and ReMoDiffuse[37] to train the retrieval and generation models respectively.
We use OpenAI’s GPT-3.5-turbo-instruct for its efficiency in executing specific instructions and providing direct answers. Our prompting strategy allows a maximum of 256 tokens for both the prompt and the generation. The completions API, with default parameters, is used to generate the desired part-specific descriptions.
For part-specific retrieval models ($MoRAG_{p}$), a batch size of 32 is used, with other hyperparameters following the configuration in TMR [23]. The MoRAG-Diffuse model adopts settings similar to those of ReMoDiffuse [37] for HumanML3D [12]. It is trained on an NVIDIA GeForce RTX 2080 Ti with a batch size of 64, leveraging pre-trained weights from ReMoDiffuse [37] for 50k steps.
@@ 4.2 Results
5 Discussion
Challenges in Motion Retrieval Metrics: To evaluate text-to-motion retrieval methods such as TMR[23] and TMR++[5], the similarity between the text corresponding to the retrieved motion sample and the input text is computed. However, for our approach, which involves the spatial composition of motion sequences, there is no single corresponding text for the entire composed sequence. As a result, these metrics cannot be computed in our case.
Assumptions and Limitations: (1)GPT Dependency: Our approach relies on GPT-generated descriptions for motion retrieval, with accuracy dependent on the precision of these descriptions. Semantic deviations may impact retrieval quality. Feedback mechanisms or quality checks can enhance robustness and mitigate inconsistencies. (2)Limited Dataset: A constrained dataset size may lead to suboptimal retrievals, particularly for complexinputs. Expanding the dataset or integrating external data can improve retrieval precision and diversity. (3)Composition Alignment: Trimming retrieved motions to the shortest sequence for alignment can introduce minor semantic or temporal inconsistencies. Incorporating a coherence check can ensure better fidelity and alignment with the input text.
Future scope: Future work could extend our approach to other architecture-based generative models. Incorporating additional body part information, such as finger, head, and lip movements from respective part-specific databases, would enhance the realism of generated samples and better handle more complex language descriptions.
@@ 6 Conclusion
Appendix
Figure 7: Spatial Motion Composition: Illustration of spatial motion composition using retrieved part-specific motion samples for the text: ”A person is standing on one leg and raising both hands”
7 Dataset
8 Implementation Details
We use OpenAI’s GPT-3.5-turbo-instruct for its efficiency in executing specific instructions and providing direct answers. Our prompting strategy allows a maximum of 256 tokens for both the prompt and the generation. The completions API, with default parameters, is used to generate the desired part-specific descriptions.
9 Prompt strategy
Table  3 presents the LLM outputs for various text descriptions generated using our prompt function, demonstrating its effectiveness in motion retrieval. We observe that the generated descriptions are often correlated with other body parts, which led us to train part-specific retrieval models on full-body motion sequences.
9.1 Spatial Composition
Fig. 7 illustrates the composition workflow for combining part-specific motions, $R_{part}$, retrieved from corresponding part-specific databases, $\mathcal{D}_{part}$, using the LLM-generated part-specific descriptions, textpart, where $part\in\{torso,hands,legs\}$.
9.2 Significance of ”position”
To train independent part-specific retrieval models, $MoRAG_{p}$, for $p\in\{torso,hands,legs\}$, it is essential to obtain movement information specific to each part. However, since the framework includes a composition step that combines the retrieved motions into a single motion sequence, relying solely on movement information is insufficient. The composition also requires positional information. For example, as shown in Fig. 8 for the text text = ”A person is swimming”, explicitly prompting for the positional information allows the leg description to reflect its relative position to the ground, thereby retrieving the correct motion sample. The global orientation of the leg corresponding retrieved sample is important as it will determine the global orientation of the composed motion sequence.
9.3 Issue with left/right retrieval strategy
In MoRAG, we avoided the strategy of retrieving the left and right hands and legs separately, as their movements can become asynchronous when composed into a single motion sequence. This issue is illustrated in Fig. 9, which demonstrates asynchronous motion composition for the input text ”A person is clapping his hands”.
10 Qualitative Analysis
We present qualitative analysis across three key aspects: (1) Generalizability, (2) Diversity, and (3) Zero-shot performance, for both MoRAG (retrieval model) and MoRAG-Diffuse (generative model). We also provide baseline comparisons for generalizability and zero-shot capabilities of MoRAG against TMR++[5], a state-of-the-art motion retrieval model, and ReMoDiffuse[37], which serves as the basis for MoRAG-Diffuse. Related videos can be found in project page.
10.1 Generalizability
The incorporation of part-specific descriptions texttorso, texthands and textlegs, generated by LLMs enables MoRAG to construct motion samples that accurately capture subtle variations in text, allowing for finer detail representation. By utilizing these detailed descriptions, the model retrieves more precise and nuanced motions, reflecting minor differences in body part movements or positions within the input text. This results in more realistic and contextually appropriate motion sequences. Example videos can be found in project page.
Conditioning these samples in the motion generation pipeline improves the model’s comprehension of the language space, allowing it to effectively capture variations in text descriptions. This improvement is reflected in the Multi-Modal Distance metric. (Tab. 2) Example videos can be found in project page.
10.2 Diversity
MoRAG constructs diverse set of motion samples for a given input text text by utilizing both, (1) LLMs’ ability to generate diverse text descriptions for a prompt and (2) various combinations of retrieved part-specific motion samples. Examples of these diverse samples can be found in project page.
The diverse samples produced by MoRAG improve the diversity of MoRAG-Diffuse, as indicated by the Diversity metric in Tab. 2. Related videos can be found in project page.
10.3 Zero-Shot Performance
Our approach facilitates the construction of motion sequences for unseen text phrases (zero-shot). This capability arises from two key aspects of the MoRAG framework: (1) it utilizes LLM-generated descriptions rather than the input text directly, and (2) it employs part-wise spatial composition. While the input text may be novel, the LLM-generated part-specific descriptions are not entirely unknown; relevant samples exist where the desired action is performed by specific body parts. Our method retrieves such samples independently for each body part and integrates them to construct motion sequences for unseen text descriptions. Example videos can be found at in project page.
Conditioning the samples constructed by MoRAG facilitates the generation of motion sequences for previously unseen text phrases in MoRAG-Diffuse. Example videos demonstrating this capability are available in project page.
Figure 8: Position Significance: Impact of positional information in leg descriptions on motion retrieval accuracy and global orientation consistency in composed sequences for the text: ”A person is swimming”
Figure 9: Asynchronous motion caused by separate retrieval of left and right parts: Illustration of asynchronous motion composition resulting from separate left and right hand retrievals for the text: ”A person is clapping his hands.”
text LLM Output MoRAG part-specific retrieval
texttorso texthands textlegs Rtorso Rhands Rlegs
A person is standing and raising both hands The person’s torso is upright and still, while standing tall and balanced. The person’s hands are being lifted up towards the sky, using their arms to extend upwards. The legs are steady and stable, acting as a strong foundation to support the body as the arms raise up. a person uses their hands to clap (#002573) a person raises his hands above his head. (#002315) a person raises his hands above his head. (#002315)
A person is standing and raising single hand The person’s torso is upright and still, while their arm raises up. One hand is lifted up from the side of the body, extended upwards and reaching toward the ceiling. The legs are supporting the person’s weight, standing firmly on the ground. (person stands still and lifts right hand to face and mouth area#000813) a person raises his right arm and then lowers it. (#001179) a figure claps around shoulder height (#007523)
A person is standing on one leg and raising both hands The person’s upper body is straight and centered while standing on one leg. Both hands are lifted above the head, reaching towards the sky. One leg is holding the person’s weight while the other is slightly lifted off the ground, balancing the body. a person grabs their right foot and places it on their left thigh, and balances on one foot and then does the same with the other foot. (#006123) raising and lowering arms. (#011583) a person balances on their left leg and then their right. (#009917)
A person is standing on one leg and raising single hand The person’s body is upright and balanced on one leg, with the other leg lifted off the ground. One hand is raised up in the air, reaching toward the ceiling or sky. The standing leg is firmly planted on the ground, while the other leg is lifted up and may be slightly bent or straight. the person raises their left foot up to their knee and then kicks their foot out, then returns their foot to their knee. (#004012) a person raises his right arm and then lowers it. (#001179) person balances on left leg then with arms up high has arms fully extended keeping balance on left leg (#004180)
A person is running with their arms crossed. The person’s upper body is straight and upright, while their chest and shoulders may be slightly forward as they run. The hands are crossed over the chest, alternating in front of the body as the person runs. The legs are moving back and forth in a rhythmic motion, propelling the person forward as they run. a person jogs forward with arms moving to his side. (#002755) a figure stands in place, crossing its arms (#003973) a person slowly jogs to the left to right and the jogs back into place (#004273)
A person is running with their arms stretched out to the sides. The person’s torso is upright and slightly leaning forward as they jog, with their chest and shoulders in a relaxed position. The person’s hands are held out to the sides at shoulder level, moving rhythmically with each step. The legs are moving in a bouncing motion, alternating between the left and right sides as the person jogs forward. jogging forward in medium pace. (#000972) the person is flying like a airplane. (#006433) a person slowly jogs to the left to right and the jogs back into place (#004273)
A person is walking with their arms circling around. As the person runs, their torso is upright and facing forward, with their chest and stomach comfortably relaxed. The person is moving their arms back and forth in a swinging motion, with their arms stretched out to the sides. The legs are taking turns lifting off the ground and propelling the person forward in a steady, rhythmic motion. running from side to side. (#014305) this person moves both arms out to his sides in a large circular motion then walks forward. (#005375) a person doing a casual walk (#004183)
A person is walking with their arms crossed. The person’s upper body remains upright while they walk with their arms crossed, with their spine and shoulders in a stable position. The person’s hands are placed on opposite shoulders, crossing in front of their torso as they walk. The legs are alternately lifting and stepping forward, propelling the person forward while they walk. a person walks on uneven ground whilst holding on to handrail. (#001029) a person crosses their arms in front of their chest, then drops them back at their sides. (#001014) a person walks forward while holding out their arms for balance (#002087)
Table 3: Prompt Examples : We illustrate the LLM-generated part-specific outputs for text descriptions alongside their corresponding top-1 retrieval results to demonstrate the effectiveness of our prompt strategy. The HumanML3D  [12] ID for the retrieved motions is indicated with the # symbol.
11 Metrics
For quantitative evaluations, we adopt the performance metrics used in ReMoDiffuse[37], which include R Precision, Frechet Inception Distance (FID), Multi-Modal Distance, Diversity and Multimodality. For R Precision and MultiModality, higher scores indicate superior performance. Conversely, lower scores are preferred for FID and Multi-Modal Distance. For Diversity, performance improves as the score more closely aligns with the real motions.
(1) R Precision evaluates how well the generated motion sequences semantically match the text descriptions. We calculate Top-1, Top-2, and Top-3 accuracy by measuring the Euclidean distance between each motion sequence and 32 text descriptions (one ground truth and 31 randomly selected descriptions). (2) FID calculates the distance between features extracted from real and generated motion sequences. (3) Multi-modal distance (MM Dist for short) measures the average Euclidean distance between the feature vectors of text and generated motions. (4) Diversity measures the variability and richness of the generated sequences. Average euclidean distance is calculated between the two-equal sized subsets which are randomly sampled from the generated motions from all test texts. (5) Multimodality measures the diversity of generated motions for a given text. We generate 10 pairs of motions for each text and compute average distance between the each pair feature vectors.
11.1 Challenges in applying motion retrieval metrics to spatially composed sequences
To evaluate text-to-motion retrieval methods such as TMR[23] and TMR++[5], the similarity between the text corresponding to the retrieved motion sample and the input text is computed. However, for our approach, which involves the spatial composition of motion sequences, there is no single corresponding text for the entire composed sequence. As a result, these metrics cannot be computed in our case.